OpenStudy (anonymous):

find the derivative of : y=sin^3 x tan 4x

OpenStudy (anonymous):

product rule and chain rule for this one. use \[(fg)'=f'g+g'f\] with \[f(x)=\sin^3(x),f'(x)=3\sin^2(x)\cos(x), g(x) = \tan(4s),g'(x)=4\sec^2(4x)\]